# Ke Xua,b, Fen-Fen Yanga,b, Chenggui Yuanc

### v1

@@ 1. $\bf(3)$}
$\leq\mathbb{E}\Big{[}\!\sup_{0\leq r\leq t}\int_{0}^{r}\big{(}L_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}\big{)}\,\mathrm{d}s\Big{]}+\int_{0}^{t}\mathbb{E}[\sup_{0\leq r\leq s}(K_{1}+K_{2}+K_{3})|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}$
$\quad+\sup_{0\leq r\leq s}(L_{2}+K_{1}+K_{2}+K_{3})\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]{\rm{d}}s$
$\leq\int_{0}^{t}\mathbb{E}[\sup_{0\leq r\leq s}(L_{1}+K_{1}+K_{2}+K_{3})|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}(L_{2}+K_{1}+K_{2}+K_{3})\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]{\rm{d}}s$
$\leq\hat{C}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s,$
where $\hat{C}:=\hat{C}(L_{1},L_{2},K_{1},K_{2},K_{3})$.
Note that for any $s\geq 0$,
$\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}=\sup_{0\leq r\leq s}\left(e^{\lambda r}\cdot e^{-\lambda r}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\right)\leq e^{\lambda s}\sup_{0\leq r\leq s}e^{-\lambda r}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}.$ (2.4)
For convenience, we denote
$\phi(t):=\mathbb{E}\left[\sup_{0\leq s\leq t}e^{-\lambda s}\left|Y_{s}^{\mu}-Y_{s}^{\nu}\right|^{2}\right].$
Taking expectations to (2.4), we obtain
$\mathbb{E}\left[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\right]\leq e^{\lambda s}\,\phi(s).$
Similarly,
$\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\leq e^{\lambda s}\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r}).$
Hence, according to (2.3) we have
$\phi(t)$ $\leq\mathbb{E}\Big{[}\sup_{0\leq s\leq t}\big{|}Y_{s}^{\mu}-Y_{s}^{\nu}\big{|}^{2}\Big{]}$
$\leq\hat{C}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s$
$\leq\hat{C}\int_{0}^{t}\left(e^{\lambda s}\,\phi(s)+e^{\lambda s}\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s$
$\leq\hat{C}e^{\lambda T}\int_{0}^{t}\left(\,\phi(s)+\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s.$
By using Gronwall’s inequality, we have
$\phi(t)$ $\leq\hat{C}e^{\lambda T}\int_{0}^{t}\mathbb{E}[\sup_{0\leq r\leq s}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]e^{\hat{C}e^{\lambda T}(t-s)}{\rm{d}}s$
$\leq\hat{C}e^{\lambda T}\cdot\mathbb{E}[\sup_{0\leq r\leq t}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\int_{0}^{t}e^{\hat{C}e^{\lambda T}(t-s)}{\rm{d}}s$
$=(e^{\hat{C}e^{\lambda T}t}-1)\mathbb{E}[\sup_{0\leq r\leq t}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})].$
@@
$\phi(t)$ $=\mathbb{E}[\sup_{0\leq s\leq t}e^{-\lambda s}\left|Y_{s}^{\mu}-Y_{s}^{\nu}\right|^{2}]$
$\leq(e^{\hat{C}e^{\lambda T}t}-1)\cdot\mathbb{E}[\sup_{0\leq r\leq t}e^{-\lambda r}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]$
$\leq(e^{\hat{C}e^{\lambda T}t}-1)\cdot\mathbb{E}[\sup_{t\in[0,T]}e^{-\lambda t}|X_{t}-\tilde{X}_{t}|^{2}].$
Combining the definition of the metric on space $(\mathscr{Y}_{T}(2),\bm{\rho}_{2,\infty})$, it holds that
@@
$\leq(e^{\hat{C}e^{\lambda T}t}-1)^{1/2}\bigg{[}\mathbb{E}\sup_{t\in[0,T]}e^{-\lambda t}|X_{t}-\tilde{X}_{t}|^{2}\bigg{]}^{1/2}$
@@
Taking $\lambda=ln{(\frac{ln2}{2T\hat{C}})}^{\frac{1}{T}},$ then we have $0<(e^{\hat{C}e^{\lambda T}t}-1)^{1/2}<1$.
By the contraction mapping principle, the mapping
@@ 1. $\bf(4)$}
    Proving the second-order moment boundedness
Apply Itô’s formula to $|X_{t}|^{2}$. Using Assumption (A1)
$2\langle x,b(x,\mu)\rangle+|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2}+\int_{Z}|c(x,\mu)z|^{2}\nu(dz)\leq L(1+|x|^{2}+\mathcal{W}_{2}^{2}(\mu,\delta_{0})).$
Taking expectation and noting $\mathbb{E}[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}(X_{s}),\delta_{0})]\leq\mathbb{E}|X_{s}|^{2}$, we derive
@@
By Gronwall’s inequality, we obtain that
@@
This establishes existence, uniqueness, and the second-order moment boundedness. ∎
Remark 2.2.
Similar with Tran et al.[41]. It follows from Assumption (A6) that,
$\leq\gamma_{1}|x|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mu,\delta_{0})+\eta,$
Proposition 2.3.
[…]
where $\gamma=\gamma_{1}+\gamma_{2}$.
Proof.
We hereby follow the methodology introduced in [41, Proposition 2.3 ].
$\sup_{t\in[0,T]}\mathbb{E}[|X_{t}|^{p}]\leq C(T,p).$ (2.6)
$\sup_{t\in[0,T]}\mathbb{E}\big{[}\,|X_{t}|^{q}\,\big{]}\leq C(T,q).$ (2.7)
Let $\lambda\in\mathbb{R}$ and consider any even integer $p\in[2,\,p_{0}]$. By applying Itô’s formula to the process $e^{-\lambda t}|X_{t}|^{p}$, we obtain
@@
$\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$
@@
Next, by repeatedly applying the binomial theorem, using Assumption (A5), the identity
$\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$
and Young’s inequality
@@
$\sum_{j=0}^{i}\binom{i}{j}a^{j}=(1+a)^{i},\quad\sum_{j=0}^{i}\binom{i}{j}ja^{j}=ia(1+a)^{i-1}$
valid for any $a\in\mathbb{R}$, we obtain,
@@
$\quad+{L_{3}^{2}}\Big{(}1+|X_{s}|+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big{)}^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg{(}(i-\frac{j}{2}-1)(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{2}|X_{s}|^{p-4}$
$\quad+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{k}|X_{s}|^{p-2-k}\Bigg{)}$
@@
$\quad+\sum_{j=0}^{i}|z|^{2i-j}Q_{p}\big{(}p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\big{)}$
$where\quad 2i-j\geq 2\quad\text{and}\quad$
@@
Combined with the identity
@@
which holds for any $a\in\mathbb{R}$, this yields,
@@
By choosing $\lambda=\gamma_{1}p$ and applying Assumption (A6), Remark 2.2, together with the identity
@@
Thus, by the principle of induction, we have established (2.6).
@@
First, by applying (2.8) with $\lambda=2\gamma$ and $p=2$, and using Assumption (A6), we obtain,
$e^{-2\gamma t}|X_{t}|^{2}=|x_{0}|^{2}+\int_{0}^{t}e^{-2\gamma s}\bigg{[}-2\gamma|X_{s}|^{2}+2\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\bigg{]}{{\rm{d}}}s$
$\quad+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W^{0}_{s}\rangle\Bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\lambda s}\Big{(}|X_{s-}+c(X_{s-},\mathcal{L}^{1}_{X_{s}-})z|^{2}-|X_{s-}|^{2}\Big{)}\tilde{N}({{\rm{d}}}s,{{\rm{d}}}z)$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\gamma s}|c(X_{s},\mathcal{L}^{1}_{X_{s}})z|^{2}\nu({{\rm{d}}}z){{\rm{d}}}s$
$\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}-\gamma|X_{s}|^{2}+\langle X_{s},b(X_{s},\mathcal{L}^{1}_{X_{s}})\rangle+\frac{1}{2}|\sigma(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}+\frac{1}{2}|\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}$
$\quad+\frac{1}{2}|c(X_{s},\mathcal{L}^{1}_{X_{s}})|^{2}\int_{\mathbb{R}^{d}_{0}}|z|^{2}\nu({{\rm{d}}}z)\bigg{]}{{\rm{d}}}s$
$\quad+2\int_{0}^{t}e^{-2\gamma s}\bigg{[}\langle X_{s},\sigma(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}\rangle+\langle X_{s},\sigma^{0}(X_{s},\mathcal{L}^{1}_{X_{s}}){{\rm{d}}}W_{s}^{0}\rangle\bigg{]}$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}e^{-2\gamma s}\Big{(}|X_{s-}+c(X_{s-},\mathcal{L}^{1}_{X_{s}-})z|^{2}-|X_{s-}|^{2}\Big{)}\tilde{N}({{\rm{d}}}s,{{\rm{d}}}z)$
$\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\left(-\gamma_{2}|X_{s}|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})+\eta\right){{\rm{d}}}s$
@@
From the identity
$\mathcal{W}_{2}^{2}\big{(}\mathcal{L}^{1}_{X_{s}},\delta_{0}\big{)}=\mathbb{E}\big{[}\,|X_{s}|^{2}\,\big{|}\,\mathcal{F}^{0}\big{]},$
and by applying the estimate (2.6), we deduce
@@
$\mathbb{E}[|X_{t}|^{2}]\leq\begin{cases}\big{(}|x_{0}|^{2}+\tfrac{\eta}{\gamma}\big{)}e^{2\gamma t}-\tfrac{\eta}{\gamma},&\gamma\neq 0,\\[3.99994pt] |x_{0}|^{2}+2\eta t,&\gamma=0,\end{cases}$
so that (2.5) holds in the case $p=2$. Assume now, as an induction hypothesis, that (2.5) is valid for every even integer $q\in[2,\,p-2]$, that is,
$\mathbb{E}[|X_{t}|^{q}]\leq\begin{cases}C_{q}\big{(}1+e^{\gamma qt}\big{)},&\gamma\neq 0,\\[3.99994pt] C_{q}(1+t)^{q/2},&\gamma=0,\ \big{(}q=2\ \text{or}\ (\gamma_{2}>0,\ q\in(2,p-2])\big{)},\\[3.99994pt] C_{q}(1+t)^{q},&\gamma=0,\ \gamma_{2}=0,\ q\in(2,p-2].\end{cases}$ (2.14)
@@
Therefore, (2.5) holds for the given even $p$. By induction, it follows that (2.5) is valid for all even integers $p\in[2,\,p_{0}]$, and by Hölder’s inequality, for all real $p\in[2,\,p_{0}]$. This completes the proof.

### v2

@@ 1. $\bf(3)$}
$\quad+2\,\mathbb{E}\left[\sup_{0\leq r\leq t}\int_{0}^{r}\int_{\mathbb{R}_{0}^{d}}\Big(|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}+\big(c(Y_{s^{-}}^{\mu},\mu_{s})-c(Y_{s^{-}}^{\nu},\nu_{s})\big)\,z|^{2}-|Y_{s^{-}}^{\mu}-Y_{s^{-}}^{\nu}|^{2}\Big)\widetilde{N}(\mathrm{d}s,\mathrm{d}z)\right]$
$\leq\mathbb{E}\Big[\!\sup_{0\leq r\leq t}\int_{0}^{r}\big(L_{1}|Y_{s}^{\mu}-Y_{s}^{\nu}|^{2}+L_{2}\mathcal{W}_{2}^{2}(\mu_{s},\nu_{s})\big)\,\mathrm{d}s\Big]+\frac{3}{4}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]$
$\quad+2(K_{1}+K_{2}+K_{3})\int_{0}^{t}\left(\mathbb{E}\Big[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+\mathbb{E}\Big[\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})\Big]\right)\,\mathrm{d}s$
$\leq\frac{3}{4}\,\mathbb{E}\Big[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}\Big]+\frac{\hat{C}}{4}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s,$
where $\hat{C}:=\hat{C}(L_{1},L_{2},K_{1},K_{2},K_{3})$. Hence,
$\mathbb{E}\Big[\sup_{0\leq s\leq t}\big|Y_{s}^{\mu}-Y_{s}^{\nu}\big|^{2}\Big]$ $\leq\hat{C}\int_{0}^{t}\left(\mathbb{E}[\sup_{0\leq r\leq s}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}+\sup_{0\leq r\leq s}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]\right){\rm{d}}s$
$\leq\hat{C}t\cdot\mathbb{E}[\sup_{0\leq r\leq t}|Y_{r}^{\mu}-Y_{r}^{\nu}|^{2}]+\hat{C}t\cdot\mathbb{E}[\sup_{0\leq r\leq t}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})].$
@@
$\mathbb{E}\Big[\sup_{0\leq s\leq t}\big|Y_{s}^{\mu}-Y_{s}^{\nu}\big|^{2}\Big]$ $\leq\frac{\hat{C}t}{1-\hat{C}t}\cdot\mathbb{E}[\sup_{0\leq r\leq t}\mathcal{W}_{2}^{2}(\mu_{r},\nu_{r})]$
$\leq\frac{\hat{C}t}{1-\hat{C}t}\cdot\mathbb{E}[\sup_{0\leq r\leq t}|\eta_{r}-\tilde{\eta}_{r}|^{2}].$
@@
$)\displaystyle\bm{\rho}_{2,\infty}(\Phi(X),\Phi(\tilde{X}$)) $=\bm{\rho}_{2,\infty}(Y,\tilde{Y}):=\bigg[\mathbb{E}\sup_{0\leq r\leq t}|Y_{r}-\tilde{Y}_{r}|^{2}\bigg]^{1/2}$
$\leq\bigg(\frac{\hat{C}t}{1-\hat{C}t}\bigg)^{1/2}\bigg[\mathbb{E}\sup_{0\leq r\leq t}|X_{r}-\tilde{X}_{r}|^{2}\bigg]^{1/2}$
@@
[…]
By the contraction mapping principle on $0,t_{0}]$, the mapping
@@
We now iterate this argument to cover the whole interval $[0,T]$. Let $N:=\lceil T/t_{0}\rceil$ and for $k=1,\dots,N-1$ define the shifted time intervals $I_{k}:=[kt_{0},(k+1)t_{0}\wedge T]$. Suppose that a unique solution $X^{(k-1)}$ has been constructed on $0,kt_{0}]$. Use the terminal value $X^{(k-1)}_{kt_{0}}$ as the initial condition for the problem on the next interval $I_{k}$. Define the mapping $\Phi^{(k)}$ on the space $\mathscr{Y}_{I_{k}}(2)$ exactly as before but with initial time $kt_{0}$ and initial law induced by $X^{(k-1)}_{kt_{0}}$. The same local estimate (with the same constant $\hat{C}$) is valid on any interval of length $\leq t_{0}$, hence $\Phi^{(k)}$ is again a contraction on $\mathscr{Y}_{I_{k}}(2)$ with contraction constant $q<1$. Therefore there exists a unique fixed point $X^{(k)}$ on $I_{k}$. By constructing the solutions, we have $X^{(k)}_{kt_{0}}=X^{(k-1)}_{kt_{0}}$, so the pieces $\{X^{(k)}:k=0,\dots,N-1\}$ connect together to give a unique adapted solution $X$ on $[0,T]$ satisfying (1.1). Thus the existence on $[0,T]$ is proved.
@@ 1. $\bf(4)$}
    Proving the boundedness of the second-order moment
Apply Itô’s formula to $|X_{t}|^{2}$. Using Assumption (A1), taking expectation and noting
$\mathbb{E}[\mathcal{W}_{2}^{2}(\mathcal{L}^{1}(X_{s}),\delta_{0})]\leq\mathbb{E}|X_{s}|^{2},$
we derive
@@
By Grönwall’s inequality, we obtain
@@
This establishes existence, uniqueness, and the boundedness of the second-order moment. ∎
We will need the following Lemma later.
@@
$\mathcal{W}^{2}_{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})=\mathbb{E}[|X_{s}|^{2}\mid\mathcal{F}^{0}],$ (2.8)
@@
Proposition 2.5.
Let assumptions (A1)-(A7) hold. Based on Theorem 2.3, we can further obtain for each $p\in[2,p_{0}]$, one can find a constant $C_{p}>0$ such that for all $t\geq 0$,
where $\gamma=\gamma_{1}+\gamma_{2}$.
Proof.
We hereby follow the methodology introduced in [32, Proposition 2.3 ].
$\sup_{t\in[0,T]}\mathbb{E}[|X_{t}|^{p}]\leq C(T,p).$ (2.10)
$\sup_{t\in[0,T]}\mathbb{E}\big[\,|X_{t}|^{q}\,\big]\leq C(T,q).$ (2.11)
Let $\lambda\in\mathbb{R}$ and consider any even integer $p$. By applying Itô’s formula to the process $e^{-\lambda t}|X_{t}|^{p}$, we obtain
Next, by repeatedly applying the binomial theorem, using Assumption (A5), the identity (2.4) and Young’s inequality
@@
$\sum_{j=0}^{i}\binom{i}{j}a^{j}=(1+a)^{i}\quad\textup{and}\quad\sum_{j=0}^{i}\binom{i}{j}ja^{j}=ia(1+a)^{i-1},$
which valid for any $a\in\mathbb{R}$, we obtain,
@@
$\quad+{L_{3}^{2}}\Big(1+|X_{s}|+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\Big)^{2}\sum_{j=0}^{i}\binom{i}{j}2^{j}|z|^{2i-j}L_{3}^{2i-j-2}\Bigg((i-\frac{j}{2}-1)(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{2}$
$\quad\times|X_{s}|^{p-4}+\sum_{k=2}^{2i-j-2}\binom{2i-j-2}{k}(1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]})^{k}|X_{s}|^{p-2-k}\Bigg)$
@@
$\quad+\sum_{j=0}^{i}|z|^{2i-j}Q_{p}\big(p-2,2i-j,|X_{s}|,1+\sqrt{\mathbb{E}[|X_{s}|^{2}|\mathcal{F}^{0}]}\big),$
$\text{where}\quad 2i-j\geq 2\quad\text{and}\quad$
@@
This, together with the identity
@@
which holds for any $a\in\mathbb{R}$, yields,
@@
By choosing $\lambda=\gamma_{1}p$ and applying Assumption (A6), Remark 2.1, together with Lemma 2.4
@@
Thus, by the principle of induction, we establish (2.10).
@@
First, by applying (2.12) with $\lambda=2\gamma$ and $p=2$, and using Assumption (A6), we obtain
$e^{-2\gamma t}|X_{t}|^{2}\leq|x_{0}|^{2}+2\int_{0}^{t}e^{-2\gamma s}\left(-\gamma_{2}|X_{s}|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mathcal{L}^{1}_{X_{s}},\delta_{0})+\eta\right){{\rm{d}}}s$
@@
Due to Lemma 2.4, we have $\mathcal{W}_{2}^{2}\big(\mathcal{L}^{1}_{X_{s}},\delta_{0}\big)=\mathbb{E}\big[\,|X_{s}|^{2}\,\big|\,\mathcal{F}^{0}\big],$ and by applying the estimate (2.10), we deduce
@@
$\mathbb{E}[|X_{t}|^{2}]\leq\begin{cases}\big(|x_{0}|^{2}+\tfrac{\eta}{\gamma}\big)e^{2\gamma t}-\tfrac{\eta}{\gamma}&\gamma\neq 0,\\[3.99994pt] |x_{0}|^{2}+2\eta t&\gamma=0,\end{cases}$
so that (2.9) holds in the case $p=2$. Assume now, as an induction hypothesis, (2.9) is valid for every even integer $q\in[2,\,p-2]$, that is,
$\mathbb{E}[|X_{t}|^{q}]\leq\begin{cases}C_{q}\big(1+e^{\gamma qt}\big)&\gamma\neq 0,\\[3.99994pt] C_{q}(1+t)^{q/2}&\gamma=0,\ \big(q=2\ \text{or}\ (\gamma_{2}>0,\ q\in(2,p-2])\big),\\[3.99994pt] C_{q}(1+t)^{q}&\gamma=0,\ \gamma_{2}=0,\ q\in(2,p-2].\end{cases}$ (2.18)
@@
Therefore, (2.9) holds for the given even pp. By induction, it follows that (2.9) is valid for all even integers $p\in[2,\,p_{0}]$, and by Hölder’s inequality, for all real $p\in[2,\,p_{0}]$. This completes the proof. ∎